# Automorphism group schemes and Weyl groups of gradings

### v1

@@ Abstract.
A scheme theoretic version of the automorphism group of a grading on an algebra is presented, and the classical result that shows that, over algebraically closed fields of characteristic 00, the automorphism group of a grading is the normalizer of its diagonal group is extended, over arbitrary fields, to this scheme setting.
@@ 1. Introduction
$\operatorname{\mathrm{Aut}}(\Gamma):=\{\varphi\in\operatorname{\mathrm{Aut}}(\Gamma)\mid\forall g\in G\ \exists h\in G\text{ such that }\varphi(\mathcal{A}_{g})\subseteq\mathcal{A}_{h}\}.$
@@
Assume for a while that the ground field $\mathbb{F}$ is algebraically closed of characteristic 00, and that $\mathcal{A}$ is finite dimensional. Assume too, without loss of generality, that $G$ is finitely generated, otherwise substitute $G$ by the subgroup generated by $\operatorname{\mathrm{Supp}}(\Gamma)$. Consider the group of characters $\widehat{G}=\operatorname{\mathrm{Hom}}(G,\mathbb{F}^{\times})$ (group homomorphisms). The grading $\Gamma$ induces a homomorphism (of algebraic groups)
@@
The goal is to consider the affine scheme group versions, $\mathbf{Stab}(\Gamma)$ and $\mathbf{Aut}(\Gamma)$, of the stabilizer and the automorphism group of $\Gamma$. The former is straightforward, but the latter has some subtleties.
@@
In what follows $\mathcal{A}$ will always denote a finite-dimensional nonassociative over an arbitrary ground field $\mathbb{F}$. Unadorned tensor products will be assumed to be over $\mathbb{F}$.
@@ 2. The automorphism group scheme of a grading
Given a finite-dimensional vector space $\mathcal{V}$ over a field $\mathbb{F}$ and a decomposition $\Gamma:\mathcal{V}=\bigoplus_{s\in S}\mathcal{V}_{s}$ as a direct sum of subspaces, denote by $\mathrm{GL}(V,\Gamma)$ the subgroup of the general linear group $\mathrm{GL}(V)$ consisting of those linear automorphisms $\varphi$ of $V$ such that, for any $s\in\operatorname{\mathrm{Supp}}(\Gamma):=\{s\in S\mid\mathcal{V}_{s}\neq 0\}$, there is an element $\sigma(s)\in\operatorname{\mathrm{Supp}}(\Gamma)$ such that $\varphi(\mathcal{V}_{s})=\mathcal{V}_{\sigma(s)}$. Under these circumstances, $\sigma$ is a permutation of $\operatorname{\mathrm{Supp}}(\Gamma)$ ($\sigma\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))$) with the property that $\dim_{\mathbb{F}}\mathcal{V}_{s}=\dim_{\mathbb{F}}\mathcal{V}_{\sigma(s)}$ for all $s\in\operatorname{\mathrm{Supp}}(\Gamma)$.
The affine group scheme version $\mathbf{GL}(V,\Gamma)$ works as follows. Any algebra $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ without proper idempotents is sent to
$\mathbf{GL}(V,\Gamma)(R):=\{\varphi\in\operatorname{\mathrm{Aut}}_{R-\operatorname{\mathrm{Mod}}}(\mathcal{V}\otimes R)\mid\\ \exists\sigma\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))\ \text{such that}\ \varphi(\mathcal{V}_{s}\otimes R)=\mathcal{V}_{\sigma(s)}\otimes R\ \forall s\in\operatorname{\mathrm{Supp}}(\Gamma)\}.$
@@
$\begin{split}\mathbf{GL}(V,\Gamma)(R)&:=\{\varphi\in\operatorname{\mathrm{Aut}}_{R\textrm{-Mod}}(\mathcal{V}\otimes R)\mid\exists n\in\mathbb{N},\,\exists e_{1},\ldots,e_{n}\ \text{nonzero}\\ &\qquad\ \text{orthogonal idempotents in $R$ with}\ 1=e_{1}+\cdots+e_{n},\\ &\qquad\ \text{and }\exists\sigma_{1},\ldots,\sigma_{n}\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma)),\ \text{such that}\\ &\qquad\varphi(\mathcal{V}_{s}\otimes Re_{i})=\mathcal{V}_{\sigma_{i}(s)}\otimes Re_{i}\ \forall s\in\operatorname{\mathrm{Supp}}(\Gamma),\,\forall i=1,\ldots,n\}.\end{split}$ (2.1)
That is, $\mathbf{GL}(V,\Gamma)(R)$ consists of those $R$-linear automorphisms of $\mathcal{V}\otimes R$ for which there are nonzero orthogonal idempotents $e_{1},\ldots,e_{n}$, with $1=e_{1}+\cdots e_{n}$, such that the restriction of $\varphi$ to each $\mathcal{V}\otimes Re_{i}$ permutes the nonzero components $\mathcal{V}_{s}\otimes Re_{i}$. Note that the permutation may be different for each $i=1,\ldots,n$.
For any homomorphism $f\colon R\rightarrow S$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$, the definition of $\mathbf{GL}(V,\Gamma)(f)$ is just given by extension of scalars from $R$ to $S$ via $f$.
It is not difficult to check that $\mathbf{GL}(V,\Gamma)$ is indeed an affine group scheme (i.e., representable). For this, let $H$ be the subgroup of $\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))$ consisting of those permutations $\sigma$ with $\dim_{\mathbb{F}}\mathcal{V}_{s}=\dim_{\mathbb{F}}\mathcal{V}_{\sigma(s)}$ for any $s\in\operatorname{\mathrm{Supp}}(\Gamma)$. Write $\operatorname{\mathrm{Supp}}(\Gamma)=\{s_{1},\ldots,s_{m}\}$ and take bases $(v_{1},\ldots,v_{n_{1}})$ of $\mathcal{V}_{s_{1}}$, $(v_{n_{1}+1},\ldots,v_{n_{1}+n_{2}})$ of $\mathcal{V}_{s_{2}}$, …., $(v_{n_{1}+\cdots+n_{m-1}+1},\ldots,v_{n_{1}+\cdots+n_{m}})$ of $\mathcal{V}_{s_{m}}$. Write $I_{1}=\{1,\ldots,n_{1}\}$, $I_{2}=\{n_{1}+1,\ldots,n_{1}+n_{2}\}$, …, $I_{m}=\{n_{1}+\cdots+n_{m-1}+1,\ldots n_{1}+\cdots+n_{m}\}$. Then $\mathbf{GL}(V,\Gamma)$ is represented by the algebra
@@
$\mathcal{A}_{\sigma}=\mathbb{F}[X_{ij},Y]/\left(\det(X_{ij}Y-1,\,X_{ij}\mid i\in I_{p},\,j\in I_{q},\,s_{q}\neq\sigma(s_{p})\right).$
@@
The next theorem is the main result of this note. It shows that the situation for algebraically closed fields of characteristic 00 extends to arbitrary fields if affine group schemes are used.
@@
Let $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ be a grading by an abelian group on a a finite-dimensional algebra $\mathcal{A}$ over an arbitrary field $\mathbb{F}$. Then,
@@
For the converse, let $R$ be in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ and let $\varphi$ be an element of the centralizer $\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl{(}\mathbf{Diag}(\Gamma)\bigr{)}$. Let $R^{\prime}=RG$ be the group algebra of $G$ over $R$ and consider the ‘generic’ diagonal automorphism $\psi\in\mathbf{Diag}(\Gamma)(R^{\prime})$ given by
@@
In particular, over an algebraically closed field of characteristic 00, so that all affine group schemes are smooth, if $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ is a grading by an abelian group on a a finite-dimensional algebra $\mathcal{A}$, then we recover the result [EK13, (1.6)] that $\operatorname{\mathrm{Aut}}(\Gamma)=\mathbf{Aut}(\Gamma)(\mathbb{F})$ is $\operatorname{\mathrm{Norm}}_{\operatorname{\mathrm{Aut}}(\mathcal{A})}\bigl{(}\operatorname{\mathrm{Diag}}(\Gamma)\bigr{)}=\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})(\mathbb{F})}\bigl{(}\mathbf{Diag}(\Gamma)(\mathbb{F})\bigr{)}$.
@@
If $\mathbf{H}$ is not smooth the equality $\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})=\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl{(}\mathbf{H}(\mathbb{F})\bigr{)}$ is not valid in general, as the following example shows.
Example 2.5.
Let $\mathcal{A}=\mathbb{F}e_{1}\oplus\mathbb{F}e_{2}$ be the two-dimensional algebra over an arbitrary field of characteristic $3$ with multiplication given by $e_{1}^{2}=e_{2}$, $e_{2}^{2}=e_{1}$ and $e_{1}e_{2}=e_{2}e_{1}=0$. This is the split two-dimensional para-Hurwitz algebra (see, e.g., [Eld21]). The affine group scheme of automorphisms $\mathbf{Aut}(\mathcal{A})$ is easily seen to be the semidirect product $\boldsymbol{\mu}_{3}\rtimes\mathsf{C}_{2}$ (see, e.g., [EL21, Example 3.5]), where $\boldsymbol{\mu}_{3}$ (the affine group scheme of third roots of unity) is the diagonal group scheme of the grading $\Gamma$ by the cyclic group $C_{3}=\langle g\rangle$, with $\deg(e_{1})=g$, $\deg(e_{2})=g^{2}$, and the constant group scheme $\mathsf{C}_{2}$ is induced by the automorphism that swaps $e_{1}$ and $e_{2}$.
It is easy to check that $\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}(\boldsymbol{\mu}_{3})=\boldsymbol{\mu}_{3}=\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}(\boldsymbol{\mu}_{3})$, and hence, since the characteristic is $3$, $\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})}(\boldsymbol{\mu}_{3})(\mathbb{F})=1$. On the other hand, $\operatorname{\mathrm{Aut}}(\mathcal{A})$ is the cyclic group $C_{2}$ of order $2$, so we have
$\operatorname{\mathrm{Norm}}_{\mathbf{Aut}(\mathcal{A})(\mathbb{F})}\bigl{(}\boldsymbol{\mu}_{3}(\mathbb{F})\bigr{)}=\operatorname{\mathrm{Norm}}_{\operatorname{\mathrm{Aut}}(\mathcal{A})}(1)=\operatorname{\mathrm{Aut}}(\mathcal{A})=C_{2}\neq 1.$
@@ 3. Weyl groups
However, $\mathcal{A}_{\overline{\mathbb{Q}}}=\overline{\mathbb{Q}}1\oplus\overline{\mathbb{Q}}w\oplus\overline{\mathbb{Q}}w^{2}$, with $w=\frac{1}{\sqrt[3]{2}}u$, that satisfies $w^{3}=1$. Hence $\mathcal{A}_{\overline{\mathbb{Q}}}$ is isomorphic to $\overline{\mathbb{Q}}\times\overline{\mathbb{Q}}\times\overline{\mathbb{Q}}$. Here the automorphism determined by $\varphi(w)=w^{2}$ belongs to $\operatorname{\mathrm{Aut}}(\Gamma_{\overline{\mathbb{Q}}})$, and it actually generates it. It follows that the Weyl group $W(\Gamma_{\overline{\mathbb{Q}}})$ is the cyclic group of order $2$, and hence $\mathsf{W}(\Gamma)$ is the associated constant group scheme $\mathsf{C}_{2}$.
@@
$1\longrightarrow\boldsymbol{\mu}_{3}\longrightarrow\mathbf{Aut}(\Gamma)\longrightarrow\mathsf{C}_{2}\longrightarrow 1$ (3.3)
Since $\mathbf{Aut}(\Gamma)(\mathbb{Q})$ is trivial, it turns out that the quotient map $\mathbf{Aut}(\Gamma)\rightarrow\mathsf{C}_{2}$ is not surjective for rational points. It also follows that the short exact sequence (3.3) does not split, because the trivial group $\operatorname{\mathrm{Aut}}(\Gamma)=\mathbf{Aut}(\Gamma)(\mathbb{Q})$ is trivial, so it does not contain cyclic subgroups of order $2$.

### v2

@@ Abstract.
A scheme theoretic version of the automorphism group of a grading on an algebra is presented, and the classical result that shows that, over algebraically closed fields of characteristic 0, the automorphism group of a grading is the normalizer of its diagonal group is extended, over arbitrary fields, to this scheme setting.
@@
Dedicated to professor Leonid Makar-Limanov, on the occasion of his 80th birthday
@@ 1. Introduction
$\operatorname{\mathrm{Aut}}(\Gamma):=\{\varphi\in\operatorname{\mathrm{Aut}}(\mathcal{A})\mid\forall g\in G\ \exists h\in G\text{ such that }\varphi(\mathcal{A}_{g})\subseteq\mathcal{A}_{h}\}.$
@@
Assume for a while that the ground field $\mathbb{F}$ is algebraically closed of characteristic 0, and that $\mathcal{A}$ is finite dimensional. Assume too, without loss of generality, that $G$ is finitely generated, otherwise substitute $G$ by the subgroup generated by $\operatorname{\mathrm{Supp}}(\Gamma)$. Consider the group of characters $\widehat{G}=\operatorname{\mathrm{Hom}}(G,\mathbb{F}^{\times})$ (group homomorphisms). The grading $\Gamma$ induces a homomorphism (of algebraic groups)
@@
The goal is to consider the affine group scheme versions, $\mathbf{Stab}(\Gamma)$ and $\mathbf{Aut}(\Gamma)$, of the stabilizer and the automorphism group of $\Gamma$. The former is straightforward, but the latter has some subtleties.
@@
In what follows $\mathcal{A}$ will always denote a finite-dimensional nonassociative algebra over an arbitrary ground field $\mathbb{F}$. Unadorned tensor products will be assumed to be over $\mathbb{F}$.
@@ 2. The automorphism group scheme of a grading
Given a finite-dimensional vector space $\mathcal{V}$ over a field $\mathbb{F}$ and a decomposition $\Gamma:\mathcal{V}=\bigoplus_{s\in S}\mathcal{V}_{s}$ as a direct sum of subspaces, denote by $\mathrm{GL}(\mathcal{V},\Gamma)$ the subgroup of the general linear group $\mathrm{GL}(\mathcal{V})$ consisting of those linear automorphisms $\varphi$ of $\mathcal{V}$ such that, for any $s\in\operatorname{\mathrm{Supp}}(\Gamma):=\{s\in S\mid\mathcal{V}_{s}\neq 0\}$, there is an element $\sigma(s)\in\operatorname{\mathrm{Supp}}(\Gamma)$ such that $\varphi(\mathcal{V}_{s})=\mathcal{V}_{\sigma(s)}$. Under these circumstances, $\sigma$ is a permutation of $\operatorname{\mathrm{Supp}}(\Gamma)$ ($\sigma\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))$) with the property that $\dim_{\mathbb{F}}\mathcal{V}_{s}=\dim_{\mathbb{F}}\mathcal{V}_{\sigma(s)}$ for all $s\in\operatorname{\mathrm{Supp}}(\Gamma)$.
The affine group scheme version $\mathbf{GL}(\mathcal{V},\Gamma)$ works as follows. Any algebra $R$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ without proper idempotents is sent to
$\mathbf{GL}(\mathcal{V},\Gamma)(R):=\{\varphi\in\operatorname{\mathrm{Aut}}_{R-\operatorname{\mathrm{Mod}}}(\mathcal{V}\otimes R)\mid\\ \exists\sigma\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))\ \text{such that}\ \varphi(\mathcal{V}_{s}\otimes R)=\mathcal{V}_{\sigma(s)}\otimes R\ \forall s\in\operatorname{\mathrm{Supp}}(\Gamma)\}.$
@@
$\begin{split}\mathbf{GL}(\mathcal{V},\Gamma)(R)&:=\{\varphi\in\operatorname{\mathrm{Aut}}_{R\textrm{-Mod}}(\mathcal{V}\otimes R)\mid\exists n\in\mathbb{N},\,\exists e_{1},\ldots,e_{n}\ \text{nonzero}\\ &\qquad\ \text{orthogonal idempotents in $R$ with}\ 1=e_{1}+\cdots+e_{n},\\ &\qquad\ \text{and }\exists\sigma_{1},\ldots,\sigma_{n}\in\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma)),\ \text{such that}\\ &\qquad\varphi(\mathcal{V}_{s}\otimes Re_{i})=\mathcal{V}_{\sigma_{i}(s)}\otimes Re_{i}\ \forall s\in\operatorname{\mathrm{Supp}}(\Gamma),\,\forall i=1,\ldots,n\}.\end{split}$ (2.1)
That is, $\mathbf{GL}(\mathcal{V},\Gamma)(R)$ consists of those $R$-linear automorphisms of $\mathcal{V}\otimes R$ for which there are nonzero orthogonal idempotents $e_{1},\ldots,e_{n}$, with $1=e_{1}+\cdots+e_{n}$, such that the restriction of $\varphi$ to each $\mathcal{V}\otimes Re_{i}$ permutes the nonzero components $\mathcal{V}_{s}\otimes Re_{i}$. Note that the permutation may be different for each $i=1,\ldots,n$.
For any homomorphism $f\colon R\rightarrow S$ in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$, the definition of $\mathbf{GL}(\mathcal{V},\Gamma)(f)$ is just given by extension of scalars from $R$ to $S$ via $f$.
It is not difficult to check that $\mathbf{GL}(\mathcal{V},\Gamma)$ is indeed an affine group scheme (i.e., representable). For this, let $H$ be the subgroup of $\operatorname{\mathrm{Sym}}(\operatorname{\mathrm{Supp}}(\Gamma))$ consisting of those permutations $\sigma$ with $\dim_{\mathbb{F}}\mathcal{V}_{s}=\dim_{\mathbb{F}}\mathcal{V}_{\sigma(s)}$ for any $s\in\operatorname{\mathrm{Supp}}(\Gamma)$. Write $\operatorname{\mathrm{Supp}}(\Gamma)=\{s_{1},\ldots,s_{m}\}$ and take bases $(v_{1},\ldots,v_{n_{1}})$ of $\mathcal{V}_{s_{1}}$, $(v_{n_{1}+1},\ldots,v_{n_{1}+n_{2}})$ of $\mathcal{V}_{s_{2}}$, …., $(v_{n_{1}+\cdots+n_{m-1}+1},\ldots,v_{n_{1}+\cdots+n_{m}})$ of $\mathcal{V}_{s_{m}}$. Write $I_{1}=\{1,\ldots,n_{1}\}$, $I_{2}=\{n_{1}+1,\ldots,n_{1}+n_{2}\}$, …, $I_{m}=\{n_{1}+\cdots+n_{m-1}+1,\ldots n_{1}+\cdots+n_{m}\}$. Then $\mathbf{GL}(\mathcal{V},\Gamma)$ is represented by the algebra
@@
$\mathcal{A}_{\sigma}=\mathbb{F}[X_{ij},Y]/\left(\det(X_{ij})Y-1,\,X_{ij}\mid i\in I_{p},\,j\in I_{q},\,s_{q}\neq\sigma(s_{p})\right).$
@@
The next theorem is the main result of this note. It shows that the situation for algebraically closed fields of characteristic 0 extends to arbitrary fields if affine group schemes are used.
@@
Let $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ be a grading by an abelian group on a finite-dimensional algebra $\mathcal{A}$ over an arbitrary field $\mathbb{F}$. Then,
@@
For the converse, let $R$ be in $\operatorname{\mathrm{Alg_{\mathbb{F}}}}$ and let $\varphi$ be an element of the centralizer $\mathbf{Cent}_{\mathbf{Aut}(\mathcal{A})}\bigl(\mathbf{Diag}(\Gamma)\bigr)(R)$. Let $R^{\prime}=RG$ be the group algebra of $G$ over $R$ and consider the ‘generic’ diagonal automorphism $\psi\in\mathbf{Diag}(\Gamma)(R^{\prime})$ given by
@@
In particular, over an algebraically closed field of characteristic 0, so that all affine group schemes are smooth, if $\Gamma:\mathcal{A}=\bigoplus_{g\in G}\mathcal{A}_{g}$ is a grading by an abelian group on a a finite-dimensional algebra $\mathcal{A}$, then we recover the result [EK13, (1.6)] that $\operatorname{\mathrm{Aut}}(\Gamma)=\mathbf{Aut}(\Gamma)(\mathbb{F})$ is $\operatorname{\mathrm{Norm}}_{\operatorname{\mathrm{Aut}}(\mathcal{A})}\bigl(\operatorname{\mathrm{Diag}}(\Gamma)\bigr)=\mathbf{Norm}_{\mathbf{Aut}(\mathcal{A})(\mathbb{F})}\bigl(\mathbf{Diag}(\Gamma)(\mathbb{F})\bigr)$.
@@
If $\mathbf{H}$ is not smooth the equality $\mathbf{Norm}_{\mathbf{G}}(\mathbf{H})(\mathbb{F})=\operatorname{\mathrm{Norm}}_{\mathbf{G}(\mathbb{F})}\bigl(\mathbf{H}(\mathbb{F})\bigr)$ is not valid in general.
@@ 3. Weyl groups
[…]
@@
$1\longrightarrow\boldsymbol{\mu}_{3}\longrightarrow\mathbf{Aut}(\Gamma)\longrightarrow\mathsf{C}_{2}\longrightarrow 1.$ (3.3)
Since $\mathbf{Aut}(\Gamma)(\mathbb{Q})$ is trivial, it turns out that the quotient map $\mathbf{Aut}(\Gamma)\rightarrow\mathsf{C}_{2}$ is not surjective for rational points. It also follows that the short exact sequence (3.3) does not split, because the group $\operatorname{\mathrm{Aut}}(\Gamma)=\mathbf{Aut}(\Gamma)(\mathbb{Q})$ is trivial, so it does not contain cyclic subgroups of order $2$.